For this question I keep coming up with 1/7, but it's marked incorrect. Can you please help me figure out what I'm not doing correctly?
The log base is 4. Find the value of x in the following equation.
log_4 x+7=log_4 50+log_4 x

Answer by ankor@dixie-net.com(22740)

he log base is 4. Find the value of x in the following equation.
log_4 x+7=log_4 50+log_4 x
:
I can't see why your answer is incorrect
:
Since you have base 4 in all the terms, it does not matter what base you use:
:
log(x+7) = log(50*x); addition of logs is the same as multiplying
Therefore:
x + 7 = 50x
7 = 50x - x
49x = 7
x = 7/49
x = 1/7
:
Check on a calc:
ln(50/7) = 1.996
:
ln(50) + ln(1/7) = 1.996, equality reigns